OpenStudy (anonymous):

George and Carmen went on a bicycle trip. They took a bus to their starting point, and then biked the rest. They traveled 275 kilometers in total, and they biked 55 kilometers more than they were bussed. Let x = kilometers traveled by bike and y = kilometers traveled by bus. Find how many kilometers they traveled via bike.

OpenStudy (anonymous):

X + Y = 275 (Total distance traveled) Y + 55 = X (Bus + 55 = Bike OR They Biked 55 more than they bussed) Substitute this value of x, (Y + 55 = X) into the first equation. (Y +55 ) + y = 275 2Y + 55 = 275 2Y = 220 Y = 110 X = Y +55 X = 110+55 X = 165 X = 165, Y = 110.

OpenStudy (anonymous):

Thanks! =]